OpenStudy (anonymous):

Harold has three dogs. The Irish terrier weighs one-third as much as the golden retriever. The golden retriever weighs half as much as the Saint Bernard. If Harold's Saint Bernard weighs one hundred fifty pounds, what is the combined weight of all three dogs? two hundred lb two hundred twenty-five lb two hundred fifty lb two hundred seventy-five lb

OpenStudy (anonymous):

i like dogs

OpenStudy (anonymous):

that helps get outta here

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

ok lmao

OpenStudy (anonymous):

the golden retriever is 1/2 the size of the saint bernard so 150/2 which = 75. So the retriever is 75 pounds. The terrier is 1/3 of 75 so 75/3 = 25. The terrier is 25 pounds. Saint Bernard - 150 Golden Retriever - 75 Irish Terrier - 25 150 + 75 + 25 = 250 The answer is C.

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

right it has to be 250

OpenStudy (anonymous):

Tanks for all the help @netertareh Couldn't have done it without you.